Question 2: Fixed costs $550,000 Variable cost per procedure $70 Revenue per procedure $175 Given the information above: What volume is required to break even?. What volume is required to provide a profit of $180,000?. What is the contribution margin?
Answer #1

Break-even volume Q = fixed cost/(revenue - variable cost) = 550,000/(175-70) = 5,238.10

Volume to provide a profit of 180,000 = (fixed cost + profit)/(revenue - variable cost) = (550,000+180,000)/(170-75) = 6,952.38

Contribution margin = revenue - variable cost = 175 - 70 = 105